# cVLA: Towards Efficient Camera-Space VLAs

### v1

@@ 5.1 Action Encoding, Depth, and Domain Randomization Ablations
[…]
@@ 5.2 One-Shot Imitation Experiments
[…]
@@ 5.3 Inference-Time Strategies
[…]
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Appendix A Pose and Trajectory L1 Metric
[…]
@@
  • $N$ is the number of poses in the trajectory.
@@
  • ($\angle$($\cdot$)∠ ( ⋅ ) denotes the angle (in degrees) of the relative rotation.
@@ Appendix D Simulation and Real Experiments
For the real data evaluations, we used different subsets of the DROID dataset [12]. This data first needs to be filtered as the quality of the extrinsic calibration is very variable, something that has been noted and systemically addressed in recent work here. For our small-scale evaluation sets, we manually filtered the data using the projection of the end-effector position into the image. For our testing purposes, after manual filtering based on calibration, we selected text prompts containing the word “block” and took 160 episodes without background clutter, which made our DROID-hard subset. This subset contains images of blocks as well as a few visually confounding items. The presence of these leads to false predictions and makes it suitable to evaluate predictions of multiple trajectories. To make predictions without confounding objects, we created DROID-easy subset, which blurred out the confounding objects, see Fig. 8 for example.
@@ Appendix H Beam-Search-NMS Implementation
We propose a variant of beam-search that also does non-maximum suppression over a span of spatially contiguous tokens. This is done in the following manner: a Point $x$ is a local maximum if $p(x)\geq p(x^{\prime})\ \forall x^{\prime}\in[x-w,x+w]$. With a noisy distribution, $w$ should be larger; however, with too large $w$, we will suppress all maxima except the global one. Thus, we find that $w=100$ was sufficiently large. Our decoding procedure is beam search with $n=3$. After processing the coordinate or angle distribution with NMS and suppressing all non-maxima by setting the token log-probability to $\infty$- ∞, the next top $n$ tokens are selected.

### v2

@@ 5.1 Action Encoding, Depth, and Domain Randomization Ablations
CLEVR Objaverse Augs. Depth Objaverse SR ($\uparrow)$ CLEVR SR ($\uparrow)$
@@ 5.2 One-Shot Imitation Experiments
Train CLEVR Sim. SR ($\uparrow$) Static data traj. L1 error ($\downarrow$)
@@ 5.3 Inference-Time Strategies
Greedy search Sampling Beam search Beam search-NMS
Top-1 34.44 34.31 34.17 33.42
@@ Acknowledgments
This research was funded by the German Research Foundation (DFG) 417962828, 539134284, and 499552394 (SFB 1597 - Small Data). This work was partially funded by the Carl Zeiss Foundation with the ReScaLe project. Nick Heppert is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
@@ Appendix A Pose and Trajectory L1 Metric
$\text{mean taj. L1}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\mathbf{t}_{i}-\hat{\mathbf{t}}_{i}\|_{1}+\angle\left(\hat{R}_{i}^{-1}R_{i}\right)\right)$ (1)
@@
  • NN is the number of poses in the trajectory.
@@
  • $\angle(\cdot)$ denotes the angle (in degrees) of the relative rotation.
@@ Appendix D Simulation and Real Experiments
For the real data evaluations, we used different subsets of the DROID dataset [12]. This data first needs to be filtered as the quality of the extrinsic calibration is very variable, something that has been noted and systemically addressed in recent work here.
For our small-scale evaluation sets, we manually filtered the data using the projection of the end-effector position into the image. For our testing purposes, after manual filtering based on calibration, we selected text prompts containing the word “block” and took 160 episodes without background clutter, which made our DROID-hard subset. This subset contains images of blocks as well as a few visually confounding items. The presence of these leads to false predictions and makes it suitable to evaluate predictions of multiple trajectories. To make predictions without confounding objects, we created DROID-easy subset, which blurred out the confounding objects, see Fig. 8 for example.
@@ Appendix H Beam-Search-NMS Implementation
We propose a variant of beam-search that also does non-maximum suppression over a span of spatially contiguous tokens. This is done in the following manner: a Point $x$ is a local maximum if $p(x)\geq p(x^{\prime})\ \forall x^{\prime}\in[x-w,x+w]$. With a noisy distribution, $w$ should be larger; however, with too large $w$, we will suppress all maxima except the global one. Thus, we find that $w=100$ was sufficiently large. Our decoding procedure is beam search with $n=3$. After processing the coordinate or angle distribution with NMS and suppressing all non-maxima by setting the token log-probability to -$\infty$, the next top $n$ tokens are selected.